DETAILS LARLINALG8 3.R.027. Find JAl and A-11. 1 0 -2 А 0 3 = 2 - 5 7 6 (a) JA (b) |A-1,
Answer #1

Find tAl and All le Given A = 03 2 -57 ܘܢ @ Now find the (AL 1 0 1 IAT 0 3 2 5 7 6 w 2 C 2 0 3 57 7 .6 =1(18-14)-0(0.710) - 9Given 0 11 03 5 7 elimination Im cusing Crauss - Jordan tind inverse of the matrina method for 30 - 0.0 D 3 D 2 --5 7 6 0 -1 0 2 1 D %3 -15 26 Rz 326 32 Rz - 0 > 0 5/13 23 73 R₂ = R₂ - 2 / Rz o 1 0 26 D -243 7/13 - 23 R = R, taR3 5/3 2/3 X3 /g6 0 03 - 4s TAL 3 43 he 5 26 326 - 13 9 13 1. g6 F 13 1 13 [5 -3 13/-15 + 36 9. - 호크 338 그 13 15f 15 338 13 338 -2 (35+30 13 338/